OpenStudy (anonymous):

If y varies directly as x, find the direct variation equation for the situation. y = 4 when x = 8

OpenStudy (anonymous):

HI

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

NEED HELP

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

WHAT IS THE AEQUATION

OpenStudy (anonymous):

TYPO

OpenStudy (anonymous):

my topic is the equation is the equations

OpenStudy (anonymous):

do you mean like y=kx?

OpenStudy (phi):

y varies directly as x that means you have an equation y = k x where k is some number they want you to find k, and then write the correct equation

OpenStudy (anonymous):

YES

OpenStudy (phi):

to find k, replace the letters y and x with the numbers they gave you put the numbers into y= k x can you do that ?

OpenStudy (anonymous):

so 4=k8?

OpenStudy (phi):

now "solve for k" you divide both sides by 8 can you do that ?

OpenStudy (kitten_is_back):

y+=+%28kx%29 k being the constant of proportionality. y = 27 when x = 21 27 = k*21 27/21 = k

OpenStudy (phi):

**so 4=k8?*** \[ \frac{4}{8}= \frac{k\cdot 8}{8}\] on the right side 8/8 is 1 and 1*k is k what do you get on the left side?

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

or 1/2?

OpenStudy (anonymous):

YEP

OpenStudy (phi):

so now you can write down the "direct equation of variation" 1/2 or 0.5 mean the same thing. so the answer (i.e. the equation they want) is y = (1/2) x or y= 0.5 x

OpenStudy (anonymous):

ok thxs

OpenStudy (anonymous):

big help

OpenStudy (anonymous):

MEDAL